OpenStudy (anonymous):

Please help me! I will fan and metal!! Using the conversion factor from eV (electron volts) to Joules, determine which energy level to another would show a line at 435 nm in an emission spectrum. Use the formulas and the energy level information below to answer the questions. Needed constants: 1.00 eV=1.6x10^-19 J C= 3.00 x10^8 m/s n=6.63x10^-34 J*s Energy level values: E6: E= -0.378 eV E5: E=-0.544 eV E4: E= -0.850 eV E3: E= 1.51 eV E2: E= -3.403 V Yes, the last one only has a V.

OpenStudy (anonymous):

For anyone that may wonder in the future, The conversion factor is 1.60x10^-19J/eV The energy for the wavelength 435nm is E = hc/λ where h is Planck's constant =6.63x10^-34J So E = 6.63x10^-34*3.00x10^8/435x10^-9 = 4.57x10^-19J 4.57x10^-19J*1eV/1.60x10^-19 = 2.86eV rounded to 2.9eV This corresponds to a drop from n = 5 to n =2 (-0.5 + 3.4 = 2.9eV) Is the answer.